PHOTOELECTRON SPECTROSCOPY OF ALUMINUM DOPED BORON CLUSTERS

\underline{WEI-LI LI}, CONSTANTIN ROMANESCU, LAI-SHENG WANG, Brown University, Chemistry Department, 324 Brook Street,Providence, RI 02912, USA.

Anionic boron clusters have been shown to be planar or quasi-planar up to B_{21}^{−} from a series of combined photoelectron spectroscopy and theoretical studies. All these boron clusters consist of a peripheral ring characterized by strong twocenter-two-electron (2c-2e) B-B bonds and one or more interior atoms. The propensity for planarity is due to σ- and π-electron delocalizations throughout the molecular plane, giving rise to concepts of σ- and π-aromaticity. The quasiplanarity, on the other hand, can be mechanical in nature - the circumference of the cluster is too small to fit the inner atoms - even for doubly aromatic clusters. Two questions arise: firstly, can isoelectronic substitution by a single aluminum atom on the outer ring enhance the planarity of quasi-planar structures, and, secondly, can the interior boron atoms be replaced by aluminum? A series of aluminum isoelectronic substitution of boron clusters have been investigated ranging from B_{7}^{−} to B_{12}^{−}. Aluminum turns out to avoid the central position in the all these clusters and enhance the planarity of AlB_{6}^{−} and AlB_{11}^{−} clusters by expanding the peripheral ring.

References:

[1] C. Romanescu, A. P. Sergeeva, W. L. Li, A. I. Boldyrev and L. S. Wang, J. Am. Chem. Soc. 133 (22), 8646-8653 (2011)

[2] T. R. Galeev, C. Romanescu, W. L. Li, L. S. Wang and A. I. Boldyrev, J. Chem. Phys. 135, (8) 104301 (2011)

[3] W. L. Li, C. Romanescu, T. R. Galeev, L. S. Wang and A. I. Boldyrev, J. Phys. Chem. A 115 (38), 10391-10397